### v1

@@ 3.1 Proprioception–Depth Input Encoding
As described in Section 1, we consider an agent that uses both proprioceptive state and depth image for decision making. At time $t$, the observation is
@@
where $s^{\text{prop}}_{t}\in\mathbb{R}^{D_{p}}$ denotes the proprioceptive vector and $I^{\text{depth}}_{t}\in\mathbb{R}^{H\times W}$ denotes a first-person depth image that captures obstacles and terrain ahead of the robot. Temporal aggregation of the proprioceptive vector and the four most recent depth frames is handled by the fusion backbone described in Section 3.2.
@@
We encode each modality with a lightweight, domain-specific network and unify them in a shared latent space. The proprioceptive vector is mapped by an MLP to a compact token:
@@
The depth image is processed by a compact CNN to produce a spatial feature map:
@@
To form a single cross-modal stream, we project both embeddings to a common width $d$. The projected proprioceptive token is
@@
where $W_{p}$ and $W_{v}$ are learned projections and $N=N_{h}N_{w}$ denotes the number of spatial tokens. We then concatenate the modalities into one token sequence
@@
Finally, we add learned position indices and modality tags, followed by per-token layer normalization, to obtain
$\hat{U}_{t}=\mathrm{LN}\!\left(U_{t}+E_{\text{pos}}^{\text{spat}}+E_{\text{mod}}\right).$ (8)
@@ 3.2 State Space Modelling with Mamba
We fuse the cross-modal token sequence $\hat{U}_{t}$ with a stack of Mamba SSM layers (Fig. 2). Each layer scans the token stream with input-dependent state updates and carries a compact recurrent state across tokens and time steps, enabling efficient long-horizon modeling.
Figure 2: Schematic of a Mamba SSM block.
@@
where $\bar{A}_{t,k},\bar{B}_{t,k},\bar{C}_{t,k},\bar{D}_{t,k}$ are input-gated, token-dependent parameters produced by lightweight affine transforms of $u_{t,k}$. The scan proceeds over $k{=}1\ldots(1{+}N)$, and the initial state at each time step is carried from the previous step, $x_{t,1}\leftarrow x_{t-1,\,1+N}$. The output tokens of the layer are $Y_{t}=\{y_{t,k}\}_{k=1}^{1+N}\in\mathbb{R}^{(1+N)\times d}$. We set $h$ and $d$ equal via a learned projection when needed.
@@
We stack $L_{\text{m}}$ Mamba layers with residual connections and layer normalization:
@@
with $H^{(0)}_{t}\!=\!\hat{U}_{t}$ and $Y^{(\ell)}_{t}$ computed by (9)–(10) using inputs $H^{(\ell)}_{t}$. The selective SSM update yields near-linear time and memory in the token count (1+N)(1{+}N), in contrast to the quadratic cost of global self-attention, and naturally supports streaming over time by reusing the carried state $x_{t,1}$.
@@
To obtain a compact feature for control, we perform modality-aware pooling on the final layer outputs $H^{(L_{\text{m}})}_{t}$. Let $y^{\text{prop}}_{t}$ be the proprioceptive token and $\{y^{\text{vis}}_{t,i}\}_{i=1}^{N}$ the visual tokens from $H^{(L_{\text{m}})}_{t}$. We compute
@@ 3.3 Policy Optimization with PPO
We model locomotion as a discounted MDP $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,\gamma)$ with discount $\gamma\in(0,1)$. At time $t$, the agent receives observation $o_{t}$ and obtains a fused feature $h_{t}=f_{\text{fuse}}(o_{t})\in\mathbb{R}^{d_{h}}$ from the Mamba backbone (Sections 3.2). The objective is to maximize the expected discounted return
@@
The action is also a 12-dimensional vector that controls the change of all the joint angles. We use a Gaussian policy in the unconstrained space with squashing to actuator limits:
$\tilde{a}_{t}\sim\mathcal{N}\!\big{(}\mu_{\theta}(h_{t}),\,\operatorname{diag}(\sigma_{\theta}^{2}(h_{t}))\big{)},\qquad a_{t}=a_{\max}\tanh(\tilde{a}_{t}).$ (14)
@@
We optimize the policy with PPO using the clipped surrogate. With likelihood ratio
@@ 4 Implementation
4.1 Computing Environment
We conduct all experiments on a cloud workstation equipped with an Intel Xeon Gold 6430 CPU (128 cores, 2.1 GHz base clock) and an NVIDIA GeForce RTX 4090 GPU (24 GB). The operating system is Ubuntu 22.04. We use PyBullet for physics simulation and implement all models in Python with PyTorch 2.4.1.
4.2 Simulation Environment
We evaluate across six simulated environments (Yang et al., 2021) that vary terrain difficulty, obstacle geometry, and the presence of collectible spheres (used as sparse task bonuses):
  • Wide Obstacle (Wide Obs.): flat terrain with wide cuboid obstacles; no spheres.
  • Thin Obstacle (Thin Obs.): flat terrain with numerous thin cuboid obstacles; no spheres.
  • Wide Obstacle & Spheres (Wide Obs. & Sph.): as Wide Obs., with collectible spheres for bonus reward.
  • Thin Obstacle & Spheres (Thin Obs. & Sph.): thin obstacles plus collectible spheres on flat terrain.
  • Moving Obstacle (Moving Obs.): similar to Thin Obs., but obstacles move in random directions with low-frequency updates.
  • Mountain: rugged terrain with a goal at the summit (progress toward the goal).
Figure 3 shows representative examples. Unless noted, obstacles and spheres are randomized at episode reset and remain static; only Moving Obs. updates obstacle positions during an episode.
Figure 3: Simulated environments. Panels (a) to (f): (a) Wide Obs.; (b) Thin Obs.; (c) Wide Obs. & Sph.; (d) Thin Obs. & Sph.; (e) Moving Obs.; (f) Mountain. Obstacles and spheres are randomized at reset and remain static except in (e).
@@ 4.3 RL MDP Details
The action is a 12-dimensional vector that controls the change of all the joint angles.
@@
We use a unified reward that balances forward progress, survival, and energy efficiency, with an optional sparse collection term:
$R_{t}=\alpha_{\text{fwd}}\,R^{\text{fwd}}_{t}+\alpha_{\text{energy}}\,R^{\text{energy}}_{t}+\alpha_{\text{alive}}\,R^{\text{alive}}_{t}+K_{t}\,R^{\text{sphere}}_{t},$ (20)
@@
$R^{\text{fwd}}_{t}=\begin{cases}\langle v_{t},\ e_{x}\rangle,&\text{flat terrains},\\[2.0pt] \langle v_{t},\ u^{\text{goal}}_{t}\rangle,&\text{mountain terrain},\end{cases}$ (21)
where $v_{t}$ is the base linear velocity, $e_{x}$ is the unit vector along the $x$-axis, and $u^{\text{goal}}_{t}$ is the unit direction from the robot to the goal. The energy term penalizes excessive actuation,
@@
The per-sphere term $R^{\text{sphere}}_{t}$ is a fixed bonus awarded for each collected item.
@@ 4.5 Training Schema
[…]
Table 2: PPO training hyperparameters (shared across methods).
@@
Episode horizon 1000
@@
Samples per iteration 8192
Minibatch size 256
Optimization epochs 3
PPO clip parameter $\epsilon$ 0.2
Policy learning rate $1\times 10^{-4}$
Value learning rate $1\times 10^{-4}$
Optimizer Adam
Policy initialization Standard Gaussian
@@
To improve robustness, all methods use the same physics randomization at episode reset, with parameters sampled uniformly from the ranges in Table 3. Unless otherwise noted, the randomized values remain fixed for the entire episode. The overall LocoMamba training process, incorporating PPO, domain randomization, and curriculum learning, is illustrated in Algorithm 1.
Table 3: Domain randomization ranges (sampled at episode reset; held constant per episode).
@@
Body mass $0.8,\ 1.2]\times$ default -
@@
Motor strength $0.8,\ 1.2]\times$ default -
@@
In addition to physics randomization, we inject lightweight depth noise to simulate saturated returns and minor sensor artifacts. At each step, we sample $K\sim\mathcal{U}\{3,\dots,30\}$ pixel locations per depth frame and set their values to the sensor’s maximum range (salt-like saturation). This perturbation is applied consistently across all methods.
@@
For obstacle-based scenarios, density starts from an easier setting and linearly ramps to the target distribution over training iterations. Curriculum scheduling is identical across methods.
@@
Gradients from the PPO objective flow through the projection head, the Mamba fusion backbone, and both encoders. We use advantage normalization and gradient clipping, and update for several epochs per iteration with shuffled minibatches.
@@ 5.1 Evaluation Setup
We organize our evaluation around the following research questions (RQs):
  • RQ1 (SOTA performance). Does LocoMamba, which fuses proprioception and depth with Mamba SSM backbone, achieve better locomotion performance than SOTA, as measured by return, success rate, and collision rate?
@@
  • RQ4 (Training robustness). Does the proposed end-to-end PPO training schema improve learning stability and support multi-objective optimization?
@@
We evaluate policies using (i) mean episode return and two domain-specific metrics: (ii) progress distance, defined as the displacement in meters along the task-aligned target direction, and (iii) collisions per episode, computed over a fixed episode length of 1,000 steps. Collision checks are performed at every control step, and the collision metric is reported only for obstacle-bearing scenarios and for episodes that encounter at least one obstacle.
@@
To assess the effectiveness of the proposed LocoMamba, we conduct comparisons with the following methods:
@@
  • Transformer Proprio-Vision. Self-attention over visual tokens plus a proprio token, representing attention-based cross-modal fusion (Yang et al., 2021).
  • Transformer Vision-only. Following Transformer Proprio-Vision, we also evaluate a transformer-based model using depth-only input.
All agents are trained using the same PPO schema, curriculum, and domain randomization, and are evaluated under identical settings. Where applicable, we reuse the same proprioceptive and depth encoders, as well as the same token width, to minimize confounding factors related to representation size; the only difference lies in the fusion module.
@@
We evaluate LocoMamba and its variant, Mamba Vision-Only, which uses only vision as input, against established baselines in the Thin Obstacle environment. Figure 4 shows training curves with mean and one standard deviation across seeds. Table 4 reports means and standard deviations over seeds for episode return, collisions per episode, and progress distance.
@@
Table 4: Performance on the Thin Obstacle environment.
Model Architecture Reward Collision Times Distance Moved
@@
Transformer Vision-Only 187.45 $\pm$ 93.88 - 6.75 $\pm$ 2.90
Mamba Vision-Only 28.16 $\pm$ 34.47 - 2.92 $\pm$ 1.01
@@
From Table 4, Mamba Proprio-Vision attains the best overall performance, with higher return ($762.34\pm 156.53$), fewer collisions ($72.53\pm 79.47$), and longer progress distance ($32.41\pm 5.23$) than all baselines. Relative to Transformer Proprio-Vision, it improves return by $48.9\%$, reduces collisions by $48.9\%$, and increases distance by $30.4\%$. Compared with Proprio-Only, it yields a $423.4\%$ gain in return, an $85.1\%$ reduction in collisions, and a $428.7\%$ increase in distance. The two vision-only variants move little on average, so collision counts are not informative and are omitted. These results indicate that cross-modal fusion is essential for obstacle negotiation and that the selective state-space backbone confers a clear advantage under identical training and evaluation protocols. The findings answer RQ1 affirmatively and support the effectiveness of LocoMamba. Moreover, under the proposed end-to-end PPO training protocol with a compact state-centric reward, the agent exhibits stable learning and balanced trade-offs among progress, smoothness, and safety, thereby answering RQ4.
@@ 5.3 Ablation Studies
Ablation on the Cross-Modal Fusion Backbone
We compare fusion backbones under matched modality settings. In the proprioception–vision setting, LocoMamba outperforms Transformer Proprio-Vision on all metrics in Table 4. The return rises from $511.96\pm 247.30$ to $762.34\pm 156.53$ (a $48.9\%$ increase), the number of collisions drops from $141.83\pm 158.47$ to $72.53\pm 79.47$ (a $48.9\%$ reduction), and the progress distance increases from $24.85\pm 7.34$ to $32.41\pm 5.23$ (a $30.4\%$ increase). In the vision-only setting, both methods perform poorly relative to cross-modal policies. Mamba Vision-Only remains substantially below Transformer Vision-Only in return and distance, underscoring the importance of proprioception for egocentric perception and control. These comparisons indicate that the Mamba SSM backbone is most beneficial when fusing modalities within a single policy.
To assess efficiency, Table 5 summarizes learning-curve statistics under identical training schedules. In the proprioception–vision setting, the Mamba backbone achieves a higher final reward (737.4 compared with 714.1), a steeper early-learning slope (6.41 per epoch compared with 3.66, a $75\%$ increase), higher learning efficiency (1.44 compared with 1.28), and a larger area under the learning curve per epoch (601.7 compared with 529.8). In the vision-only setting, the Mamba backbone attains a higher final reward (676.0 compared with 577.8) and a steeper early slope (4.22 compared with 3.69), with a comparable area under the curve. Taken together, these results support RQ3: the Mamba-based backbone improves optimization dynamics under the same training budget and confers efficiency benefits that translate into faster learning and stronger final performance when cross-modal fusion is used. In addition, the learning curves in Fig. 4 show faster convergence and lower variance for the Mamba-based models, indicating more stable optimization. This further supports the efficiency advantage of the Mamba SSM backbone (RQ3) and the robustness of the proposed PPO training protocol (RQ4).
@@
Model Final Reward Early Learning Slope Learning Efficiency AUC per Epoch

### v2

@@ 3.1 Proprioception–Depth Input Encoding
As described in Section 1, an agent is considered to uses both proprioceptive state and depth information for decision making. At time $t$, the observation is
@@
where $s^{\text{prop}}_{t}\in\mathbb{R}^{D_{p}}$ denotes the proprioceptive vector and $I^{\text{depth}}_{t}\in\mathbb{R}^{H\times W}$ denotes a first-person-view depth image that captures obstacles and terrain ahead of the robot. Temporal aggregation of the proprioceptive vector and the four most recent depth frames is handled by the fusion backbone described in Section 3.2.
@@
Each modality is encoded with a lightweight, domain-specific network and unify them in a shared latent space. The proprioceptive vector is mapped by an MLP to a compact token:
@@
The depth image is processed by a compact CNN [45] to produce a sequence of spatial tokens:
@@
To form a single cross-modal stream, both embeddings are projected to a common width $d$. The projected proprioceptive token is
@@
where $W_{p}$ and $W_{v}$ are learned projections, and $N$ denotes the number of spatial tokens. Then the modalities are concatenated into one token sequence
@@
Finally, learned position indices and modality tags are added, followed by per-token layer normalization, to obtain
$\hat{U}_{t}=\mathrm{LN}\!\left(U_{t}+E_{\text{pos}}^{\text{spat}}+E_{\text{mod}}\right),$ (8)
@@ 3.2 State Space Modelling with Mamba
The cross-modal token sequence $\hat{U}_{t}$ is fused with a stack of Mamba SSM layers [38] (Fig. 2). Each layer scans the token stream with input-dependent state updates and carries a compact recurrent state across tokens and time steps, enabling efficient long-horizon modeling.
Figure 2: Schematic of a Mamba SSM Layer.
@@
where $\bar{A}_{t,k},\bar{B}_{t,k},\bar{C}_{t,k},\bar{D}_{t,k}$ are input-gated, token-dependent parameters produced by lightweight affine transforms of $u_{t,k}$. The scan proceeds over $k{=}1\ldots(1{+}N)$, and the initial state at each time step is carried from the previous step, $x_{t,1}\leftarrow x_{t-1,\,1+N}$. The output tokens of the layer are $Y_{t}=\{y_{t,k}\}_{k=1}^{1+N}\in\mathbb{R}^{(1+N)\times d}$. The token width $h$ and the hidden state dimension of the SSM $d$ are set to equal by a learned projection when needed.
@@
The $L_{\text{m}}$ Mamba layers are stacked with residual connections and layer normalization:
@@
with $H^{(0)}_{t}\!=\!\hat{U}_{t}$ and $Y^{(\ell)}_{t}$ computed by (9)–(10) using inputs $H^{(\ell)}_{t}$. The selective SSM update yields near-linear time and memory in the token count (1+N)(1{+}N), in contrast to the quadratic cost of global self-attention [33], and naturally supports streaming over time by reusing the carried state $x_{t,1}$.
@@
To obtain a compact feature for control, modality-aware pooling is performed on the final layer outputs $H^{(L_{\text{m}})}_{t}$. Let $y^{\text{prop}}_{t}$ be the proprioceptive token and $\{y^{\text{vis}}_{t,i}\}_{i=1}^{N}$ be the visual tokens from $H^{(L_{\text{m}})}_{t}$. Then they can be calculated as:
@@ 3.3 Policy Optimization with PPO
The locomotion can be modeled as a discounted MDP $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,\gamma)$ with discount $\gamma\in(0,1)$[46]. At time $t$, the agent receives observation $o_{t}$ from environment and obtains a fused feature $h_{t}=f_{\text{fuse}}(o_{t})\in\mathbb{R}^{d_{h}}$ from the Mamba backbone (Sections 3.2). The objective is to maximize the expected discounted return
@@
The action is also a 12-dimensional vector that controls the change of all the joint angles. A Gaussian policy is used in the unconstrained space with squashing to actuator limits:
$\tilde{a}_{t}\sim\mathcal{N}\!\left(\mu_{\theta}(h_{t}),\,\operatorname{diag}\!\big{(}\sigma_{\theta}^{2}(h_{t})\big{)}\right).$ (14)
$a_{t}=a_{\max}\,\tanh(\tilde{a}_{t}).$ (15)
@@
The policy is optimized with PPO [44] using the clipped surrogate. With likelihood ratio
@@ 4 Implementation
4.1 Environment Setup
All experiments are conducted on a laptop equipped with an Intel Core Ultra 7 155H CPU (22 cores, 1.4 GHz base clock) and an NVIDIA GeForce RTX 500 Ada GPU (4 GB, CUDA-12.9). The operating system is Ubuntu 22.04. The physics simulation is run with PyBullet [47] and all models are implemented in Python 3.8 with PyTorch 2.4.1.
The model is evaluated across three simulated environments that vary in terrain difficulty and obstacle dynamics:
  • Thin Obstacle: flat terrain populated with numerous thin cuboid obstacles.
  • Rugged Terrain: uneven, discontinuous ground with a maximum height of 5 cm, requiring careful placement of foothold.
  • Dynamic Obstacle: thin obstacles that move in random directions.
Fig. 3 shows representative examples. Unless noted otherwise, obstacle layouts are randomized at episode reset; only Dynamic Obstacle updates obstacle positions during an episode.
Figure 3: Simulated environments. Panels (a)–(c): (a) Thin Obstacle; (b) Thin Obstacle with Rugged Terrain; (c) Dynamic Obstacle. Obstacle layouts are randomized at reset; only (c) updates obstacle positions during an episode.
@@ 4.2 RL MDP Details
The action is a 12-dimensional vector that specifies desired joint position targets for the 12 actuated joints. A proportional–derivative (PD) controller [48] converts these targets into joint torques on the agent.
@@
A unified reward is used that balances forward progress, energy efficiency, and survival:
$R_{t}=\alpha_{\text{fwd}}\,R^{\text{fwd}}_{t}+\alpha_{\text{energy}}\,R^{\text{energy}}_{t}+\alpha_{\text{alive}}\,R^{\text{alive}}_{t},$ (21)
@@
$R^{\text{fwd}}_{t}=\langle v_{t},\ e_{x}\rangle$ (22)
where $v_{t}$ is the base linear velocity, $e_{x}$ is the unit vector along the $x$-axis. The energy term penalizes excessive actuation,
@@ 4.4 Training Schema
The model is trained with PPO using on-policy rollouts of length $T$, minibatch updates over several epochs, advantage normalization, and gradient clipping. Terrain and appearance randomization are applied at episode resets, and an obstacle-density curriculum gradually increases task difficulty. The model is evaluated periodically, repeated for 10 times, and the mean and standard deviation are reported over these runs. Table 2 lists PPO hyperparameters shared across methods. The overall LocoMamba training process, incorporating PPO, domain randomization, and curriculum learning, is illustrated in Algorithm 1.
Table 2: PPO training hyperparameters.
@@
Episode horizon (steps) 999
Samples per iteration 16,384
Minibatch size 1,024
Optimization epochs per update 3
@@
GAE parameter $\lambda$ 0.95
PPO clip parameter $\epsilon$ 0.2
Entropy coefficient 0.005
Policy learning rate $2\times 10^{-4}$
Value learning rate $2\times 10^{-4}$
Optimizer Adam
@@
To improve robustness, all methods use the same physics randomization [49] at episode reset, with parameters sampled uniformly from the ranges in Table 3. Unless otherwise noted, the randomized values remain fixed for the entire episode.
Table 3: Domain randomization ranges.
@@
Body mass $0.8,\ 1.2]\times$ default Kg
@@
Motor strength $0.8,\ 1.2]\times$ default N·m
@@
In addition to physics randomization, the lightweight depth noise is also injected to simulate saturated returns and minor sensor artifacts. $K\sim\mathcal{U}\{3,\dots,30\}$ pixel locations are sampled per depth frame and their values are set to the sensor’s maximum range (salt-like saturation). This perturbation is applied consistently across all methods.
@@
For obstacle-based scenarios, density starts from an easier setting and linearly ramps to the target distribution over training iterations [50]. Curriculum scheduling is identical across methods.
@@
Gradients from the PPO objective flow through the projection head, the Mamba fusion backbone, and both encoders. The advantage normalization and gradient clipping are used and updated for several epochs per iteration with shuffled minibatches.
@@ 5.1 Evaluation Setup
Our evaluation is organized around the following research questions (RQs):
  • RQ1 (SOTA performance). Does LocoMamba, which fuses proprioception and depth with Mamba SSM backbone, achieve better locomotion performance than SOTA, as measured by return, collision times, and distance moved?
@@
  • RQ4 (Training robustness). Does the proposed end-to-end PPO training scheme improve learning stability and performance consistency?
@@
The policies are evaluated using (i) mean episode return and two domain-specific metrics: (ii) progress distance, defined as the displacement in meters along the task-aligned target direction, and (iii) collision times, computed over the entire locomotion ending either when 3 evaluation episodes complete or the robot falls. Collision checks are performed at every control step, and the collision metric is reported only for obstacles-bearing scenarios and for episodes that encounter at least one obstacle.
@@
To assess the effectiveness of the proposed LocoMamba, comparisons are conducted with the following methods:
@@
  • Transformer Proprio-Vision. Self-attention over visual tokens plus a proprio token, representing attention-based cross-modal fusion [12].
  • Transformer Vision-only. Following Transformer Proprio-Vision, a transformer-based model using depth-only input is also evaluated.
All agents are trained using the same PPO schema, curriculum, and domain randomization, and are evaluated under identical settings. Where applicable, the same proprioceptive and depth encoders, as well as the same token width, are reused to minimize confounding factors related to representation size.
@@
LocoMamba and its variant, Mamba Vision-Only, which uses only vision as input, are evaluated against established baselines in the Thin Obstacle environment. Fig. 4 shows training curves with mean and one standard deviation across seeds. Table 4 reports means and standard deviations over seeds for episode return, collisions per episode, and progress distance.
@@
Table 4: Performance on the Thin Obstacle environment. Mean $\pm$ std over 10 runs.
Model Architecture Return Collision Times Distance Moved
@@
Transformer Vision-Only 187.45 $\pm$ 93.88 6.75 $\pm$ 2.90
Mamba Vision-Only 28.16 $\pm$ 34.47 2.92 $\pm$ 1.01
@@
From Table 4, Mamba Proprio-Vision attains the best overall performance, with higher return ($762.34\pm 156.53$), fewer collisions ($72.53\pm 79.47$), and longer progress distance ($32.41\pm 5.23$) than all baselines. Relative to Transformer Proprio-Vision, it improves return by $48.9\%$, reduces collisions by $48.9\%$, and increases distance by $30.4\%$. Compared with Proprio-Only, it yields a $423.4\%$ gain in return, an $85.1\%$ reduction in collisions, and a $428.7\%$ increase in distance. The two vision-only variants move little on average, so collision counts are not informative and are omitted. These results indicate that cross-modal fusion is essential for obstacle negotiation and that the selective state-space backbone confers a clear advantage under identical training and evaluation protocols. The findings answer RQ1 affirmatively and support the effectiveness of LocoMamba. Moreover, under the proposed end-to-end PPO training protocol with a compact state-centric reward, the agent exhibits stable learning and balanced trade-offs among task-aligned progress, energy efficiency, and safety, thereby answering RQ4.
@@ 5.3 Ablation Studies
Ablation on the Cross-Modal Fusion Backbone.
The Mamba-based cross-modal backbone is now evaluated for both effectiveness and efficiency through controlled ablations against baselines.
Effectiveness of the Mamba cross-modal backbone. The fusion backbones is evaluated under matched modality settings. In the proprioception–vision setting, LocoMamba outperforms Transformer Proprio-Vision on all metrics in Table 4. The return rises from $511.96\pm 247.30$ to $762.34\pm 156.53$ (a $48.9\%$ increase), the number of collisions drops from $141.83\pm 158.47$ to $72.53\pm 79.47$ (a $48.9\%$ reduction), and the progress distance increases from $24.85\pm 7.34$ to $32.41\pm 5.23$ (a $30.4\%$ increase). In the vision-only setting, both methods perform poorly relative to cross-modal policies. Mamba Vision-Only remains substantially below Transformer Vision-Only in return and distance, underscoring the importance of proprioception for egocentric perception and control. These comparisons indicate that the Mamba SSM backbone is most beneficial when fusing modalities within a single policy.
Efficiency of the Mamba cross-modal backbone. Table 5 summarizes learning-curve statistics under identical training schedules. In the proprioception–vision setting, the Mamba backbone achieves a higher final reward (737.4 compared with 714.1), a steeper early-learning slope (6.41 per epoch compared with 3.66, a $75\%$ increase), higher learning efficiency (1.44 compared with 1.28), and a larger area under the learning curve per epoch (601.7 compared with 529.8). In the vision-only setting, the Mamba backbone attains a higher final reward (676.0 compared with 577.8) and a steeper early slope (4.22 compared with 3.69), with a comparable area under the curve. Taken together, these results support RQ3: the Mamba-based backbone improves optimization dynamics under the same training budget and confers efficiency benefits that translate into faster learning and stronger final performance when cross-modal fusion is used. In addition, the learning curves in Fig. 4 show faster convergence and lower variance for the Mamba-based models, indicating more stable optimization. This further supports the efficiency advantage of the Mamba SSM backbone (RQ3) and the robustness of the proposed PPO training protocol (RQ4).
@@
Model Return Early Learning Slope Learning Efficiency AUC per Epoch